how do we find h0 and h1

2016 Questions 1 and 2 are based on the study described below MQ Marketing Services researched mobile technology usage among students in both 2014 and 2015. A study was undertaken in which random samples of Macquarie University students were Invited to take part in a project regarding their smartphone usage. 27 students from 2014 and 29 students from 2015 particlpated in this study. The output below was constructed using the times these students daimed to typically spend per day using smartphones. Use this output to answer the questions which follow. Smenphone Dely Unge Smanghone dely u Table 1: Summary statistics for smartphone usage (minutes per day) 2014 2015 94.759 93.68 86.129 Mean Мean Median 86.5 Median 17.586 Standard Deviation 309.245 Sample Variance 66.93 Range 17.469 Standard Deviation Sample Variance 305.138 61.58 Range 61.41 55.13 Minimum 122.06 Maximum Minimum 122.99 Maximum 2748.02 2325.47 Sum Sum 29 27 Count Count 17.525 Pooled standard deviation Table 2: Table of Excel commands and associated p-values P-value Excel command T.DIST.2T(2.402,26) T.DIST.2T(2.402,54) 0.0500 T.DIST.2T(5.077,28) 0.0500 T.DIST.2T(18.485,54) P-value Excel command 0.0238 0.0771 T.DIST.2T(1.841,26) T.DIST.2T(1.841,54) T.DIST.2T(2.005,54) T.DIST.2T(2.056,26) 0.0198 0.0713 0.0000 0.0000 3
Question 1 (7 marks) A report from MQ Marketing Services In 2012 claimed that Macquarle University students spent, on average, 78 minutes per day using smartphones at that time. MQ Marketing Services wished to determine whether the average time students were spending on smartphones had changed by 2014. Use an appropriate hypothesis test to address the following research question. Research Question: Was the average smartphone usage for a student enrolled at Macquarie University in 2014 equal to 78 minutes per day?
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

Ho: The average smartphone usage for a student enrolled at Macquire Univ in 2014 = 78 per day

H1: The average smartphone usage for a student enrolled at Macquire Univ in 2014 != 78 per day

Test statistic = (Xbar - Mu)/(Stdev/sqrt(n)) = (78-86.129)/(17.586/sqrt(27)) = -2.402

So, df = n-1 = 27-1 = 26

p-value = .0238, according to the table.

This is less than .05, we reject Ho and conclude that :

The average smartphone usage for a student enrolled at Macquire Univ in 2014 is not equal to  78 per day